Mapa mental
PROBABILITY
Conteúdo do mapa
PROBABILITY
Examples
- Dice : Two dices are rolled, what is the probability of getting 12 as a sum? Solution : No. of possible outcomes = 36 No. of favourable outcomes = 1 i.e. (6,6) P(E) = 1/36
- Card : What is the probability of getting an ace from a pack of 52 cards? Solution : No. of favourable outcomes = 4 no. of possible outcomes = 52, P(E) = 4/52 = 1/13
- Coin : When a coin is tossed, what would be the probability of appearing head? Solution : Total outcomes = 2 i.e. (H,T) Favourable outcomes = 1 probability P(E) = 1/2
Value
- Sum of probabilities of all elementary events is 1. For events A,B,C;P(A)+P(B)+P(C)=1
- Sum of probabilities of all elementary events is 1. For events A,B,C;P(A)+P(B)+P(C)=1
Theoritical Probability
- 0<P(E)<1P(E)=Number of trialsnin which the event happened/total no. of trials
Experimental Probability
- What actually happens in an experiment : P(E)=Number of outcomes favourable to E/No. of all possible outcomes of the experiment
- What actually happens in an experiment : P(E)=Number of outcomes favourable to E/No. of all possible outcomes of the experiment
Definations
- Sure or certain event when event having probability to occur as 1
- Elementary event : Event having only one outcome of the experiment
- Complementary Event : For event E, complement event, P(E)=1-P(E)
Informações
- Descrição
- Mapa Mental sobre PROBABILITY, criado por Sairindhri Roy em 09-11-2021.
- Criado por
- Sairindhri Roy
- Visualizações
- 0
- Avaliação